OpenStudy (anonymous):

find an ordered triple that represents 6y +4z if y= (2,6,8) and z=( 7,-2,6)

OpenStudy (anonymous):

6y+4z for y = 2, z = 7 6y+4z for y = 6, z = -2 6y+4z for y= 8, z =6

OpenStudy (anonymous):

so i add it or what?

OpenStudy (anonymous):

you substitute the values

OpenStudy (anonymous):

so 2x7 idk

OpenStudy (anonymous):

i think its (40,28,72) am i right?

Directrix (directrix):

@cutegirl Do these calculations from @bloopman 6y+4z for y = 2, z = 7 --> 6*2 + 4*7 = ? 6y+4z for y = 6, z = -2 6y+4z for y= 8, z =6

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Re-check the last one. @cutegirl

OpenStudy (anonymous):

its right, what wrong with it @Directrix

Directrix (directrix):

I got 62. Let me check.

Directrix (directrix):

48 + 24 = 72 okay

OpenStudy (anonymous):

thanks so much i got one more @Directrix